# The Incredible Shrinking Context… in a decompiler near you

### v1

@@ Abstract.
We compare Shrnkr to state-of-the-art decompilers, both static-analysis- and symbolic-execution-based. In a standard benchmark set, Shrnkr scales to over 99.5% of contracts (compared to $\sim$95%), covers (i.e., reaches and manages to decompile) 67% more code, and reduces key imprecision metrics by over 65%.
@@ 1. Introduction
From a technical standpoint, the problem of EVM decompilation is especially challenging. The EVM bytecode language is extremely low-level with respect to control flow, replacing all execution-control constructs with jump instructions to an address popped from the execution stack. That is, all control-flow statements (e.g., conditionals, loops, function calls, function returns) are translated into jumps to an address that is a run-time value of the low-level program. To make matters worse, the compiler performs aggressive optimization in the form of basic block merging: if the same contiguous sequence of instructions arises in several code locations (e.g., in entirely different functions, or multiple times in the same function) it is represented only once in the output binary, and control flow from any of the disparate uses of these instructions jumps to the one, merged, compiled code fragment.
The challenge of EVM decompilation, thus, is to derive a higher-level representation, including functions, calls, returns, and structured control flow, from EVM bytecode. As a program analysis challenge, it has been the domain for applying several techniques. The primary distinction is between symbolic execution approaches (Kolinko and Palkeo, 2020; Becker, 2023) and static analysis (Grech et al., 2019, 2022) approaches. Symbolic-execution-based decompilers are easier to develop, naturally produce partial results (i.e., can always produce something, rather than failing on a smart contract in its entirety), yet are often vastly incomplete, failing to even discover a significant portion of the code. In contrast, static-analysis-based decompilers typically require much heavier development effort, cover (nearly) all code, but can fail to scale or can produce imprecise output.
@@
The technical essence of Shrnkr lies in several improvements over past static-analysis-based approaches: more precise and scalable fundamental static modeling, control-flow normalization via cloning, and pre-analysis-guided elimination of spurious calls. One key novelty is responsible for the lion’s share of the benefit, as demonstrated experimentally. The improvements in the fundamental static model are due to a new kind of static context kept inside the decompiler. That is, the decompiler maintains as its current control-flow history (i.e., how the execution got to the currently-analyzed statement) a list of basic blocks that is updated under a different algorithm. The new logic, dubbed shrinking context sensitivity, aggressively shrinks the context when a likely matching call-return or chained-call pattern is observed.
@@
We begin with a brief background, introducing the elements of the EVM that we will refer to in later sections.
2. Background: EVM Smart Contracts and Decompilation
Smart contracts are small programs (typically up to around 1,000 lines of high-level code) stored on a persistent blockchain as part of its state. They are typically written in a high-level programming language, with Solidity being by far the most widely used. Due to the dominance of Solidity over all other languages (either high-level or low-level) our work considers the universe of Solidity smart contracts. At the high-level, a smart contract defines a set of external/public functions, which are its public entry points through which Externally-Owned Accounts (EOAs) or other smart contracts can interact with the contract, and a set of persistent storage variables which are part of the contract’s state on the blockchain. Code reuse is facilitated through the use of internal (a.k.a. private) functions, inheritance, and library contracts. Solidity is a statically-typed language supporting operations on a number of value types (signed and unsigned integers, bytes, boolean), dynamic-length arrays, and associative mappings, as well as complex types combining the above.
@@ 2. Background: EVM Smart Contracts and Decompilation
In addition, compiler version and settings greatly affect the produced runtime bytecode. The recent release of Solidity v0.8.0 (Solidity Team, 2020) introduced checked arithmetic and enabled the use of v2 of the ABI encoder by default, greatly increasing the number of internal functions. Notably, since the release of Solidity v0.8.13 (Solidity Team, 2022) a new compilation pipeline became stable, with plans to make it the default in Solidity v0.9.0. (The definite version does not imply a definite timeline, however.) This new Yul/viaIR pipeline involves Yul: a standardized, exportable intermediate language/representation (IR). The Yul/viaIR pipeline enables deeper optimizations and more auditable code generation than the currently default “legacy” pipeline.
@@
We define the problem of EVM bytecode decompilation (a.k.a. binary lifting) as the derivation of high-level control-flow constructs and program structure from EVM bytecode. One can view the problem as the attempt to reconstruct a high-level program [^2] Notably, none of the EVM decompilers produce code that can be re-compiled. The reason has to do with a different complication of EVM compilation, the use of the temporary memory structure. Although there have been techniques (Lagouvardos et al., 2020; Albert et al., 2023a) to reverse-engineer the majority (e.g., around 90%) of memory operations into high-level constructs (e.g., array or record copies), the techniques remain incomplete, hindering the full production of re-compilable code. This does not diminish the value of EVM decompilation: the output of state-of-the-art decompilers is typically excellent both for human consumption and for writing automated program processing tools (e.g., static analyzers (Smaragdakis et al., 2021), symbolic-execution tools (Gritti et al., 2023; Ruaro et al., 2024), or program verification engines (Grossman et al., 2017; Albert et al., 2023b, 2020)). from a low-level, stack-based intermediate representation (IR), where all control flow is represented in a continuation-passing style (CPS) form. For instance, a function call is done by pushing a continuation on the stack (the address of the basic block to return to), then pushing the function’s entry block address, and jumping. All control-flow patterns, such as in-function branching, tail calls, calls in-sequence, passing a return value of a call as an argument to another, etc., are represented as complex sequences of pushing continuations and eventually jumping to the first.
@@ 3.1. Compiler Translation
One can glimpse the low-level complexity of compiled smart contracts by considering a simple example:
1 pragma solidity 0.8.25;
2 interface IERC20 {
3 function transfer(address to, uint256 value) external returns (bool);
4 }
5 contract DecomptTest2 {
@@
7
8 function transWithFee(address token, address to, uint256 ammt, uint256 feeA, uint256 feeB)
9 external {
10 IERC20(token).transfer(to, ammt - defaultFee - feeA - feeB); // 3 private function calls!
11 }
12 function simpleTransfer(address token, address to, uint256 ammt) external {
13 IERC20(token).transfer(to, ammt);
14 }
15 }
@@
The example program contains two external functions that accept various parameters and perform fund-transfer operations. As we discuss later, the expression ammt - defaultFee - feeA - feeB on line 10 actually performs three function calls to a private function used to check subtraction for underflow.
@@ 3.2. Public Function Patterns
As public functions are not inherently supported by the EVM, high-level languages adhere to the contract Application Binary Interface (ABI) (Solidity, [n. d.]) specification, which specifies how input and output data are to be encoded when interacting with a smart contract. The ABI specifies that the first 4 bytes provided in each smart contract invocation, are called the function selector, and are used to identify the public function that is being called.
Figure 2 shows the bytecode implementing the function selector logic for our example.
0x18: JUMPDEST
0x19: PUSH0
@@
0x25: PUSH2 0x38 // push function address of transWithFee() public function
@@
These compiler-produced patterns implementing the function-selector logic have been used by past tools (Grech et al., 2019, 2022; Albert et al., 2018) to identify public function entries. However even detecting such simple patterns can have challenges. Looking at the previous code segment it is trivial for a local analysis to deduce that the EQ statement 0x24 operates on the function selector, since the input variable is loaded in the same block. EQ statement 0x2f also checks the function selector, however an inter-block analysis is required to be able to deduce this. The state-of-the-art Elipmoc (Grech et al., 2022) binary lifter requires computing the public function entries before performing its global control-flow graph analysis. In order to achieve this, it uses an approximation based on a local-only analysis that does not verify that the selector is actually used.
In Section 6 we propose a 2-phase global analysis that, among others, tackles this problem.
@@ 3.3. Private Function Patterns
The basic pattern for private function calls has been identified in past literature (Grech et al., 2019, 2022; He et al., 2023):
@@
0x12d: DUP3 // position argument in correct stack position
@@
0x10b: PUSH20 0xffffffffffffffffffffffffffffffffffffffff
0x121: AND // masks argument’s upper 12 bytes off
@@
A basic block makes a call to internal function cleanup_t_uint160 at offset 0x109, having first pushed the bytecode offset at which it wants to return after the called function’s execution completes. The return block of a function is a locally unresolved block that jumps back to the continuation block pushed by its caller.
In case of optimized code, a block performing an internal function call can also push the continuations of future calls. This pattern is typically produced in cases of calls that can be chained together (such as complex arithmetic expressions). As an example, compiling our example program using the Yul pipeline and enabling the optimizer will translate ammt - defaultFee - feeA - feeB (which implicitly involves three internal calls, to functions for checked subtraction) to the EVM bytecode in Figure 3.
@@
0x60: PUSH2 0x72 // pushes continuation for 3rd call
@@
0x66: PUSH2 0x72 // pushes continuation for 2nd call
@@
0x6e: PUSH2 0x1c7 // push checked_sub_uint256 function address
@@
0x72: JUMPDEST // continuation address for 2nd, 3rd call
0x73: PUSH2 0x1c7 // push checked_sub_uint256 function address
@@
0x77: JUMPDEST
0x78: PUSH1 0x40
0x7a: MLOAD
Figure 3. Optimized Chained Function Calls
In this optimized case, block 0x58 will set the stack so that the 3 checked sub-operations are chained. Earlier work on the Elipmoc decompiler includes a function reconstruction algorithm (Grech et al., 2022, Figure 6) that recursively infers these chained function calls. It assumes that each low-level block will map to one high-level function call. However, in optimized code, the same low-level block can be used to perform more than one high-level function call. This can be seen in the above example, which pushes the address of block 0x72 on the stack twice, implementing in this way the last two checked-subtraction operations. Such complex patterns evade past function reconstruction algorithms. In Section 5 we propose a cloning-based technique that identifies such low-level blocks with different high-level uses, and clones them, recovering precision of decompilation output.
@@ 3.4. Our Context
Elipmoc’s combination of scalability, precision, and completeness, paired with its expressive IR, have established it as a dominant lifter for EVM bytecode. Research tools for diverse program analysis applications have been implemented on top of Elipmoc. The applications include static-analysis (Brent et al., 2020; Lagouvardos et al., 2020; Kong et al., 2023; Liao et al., 2022; Smaragdakis et al., 2021; Grech et al., 2020; Munir and Reichenbach, 2023; Yang et al., 2024; Zhong et al., 2024), symbolic execution (Ruaro et al., 2024; Gritti et al., 2023), and deep learning (Zhao et al., 2023).
Shrnkr’s novel techniques mainly improve the decompiler’s context-sensitive global control-flow graph with the introduction of the shrinking context sensitivity variant described in Section 4 and its tuning via incompleteness in Section 6. Section 5 describes the introduction of a block cloning transformation step, which is performed before the global analyses and helps Shrnkr produce normalized decompilation output.
@@ 4. Shrinking Context-Sensitivity
Past work (Grech et al., 2022, 2019) has established a context-sensitive global control-flow-graph analysis as the backbone of a decompiler. That is, the decompiler abstractly simulates all possible executions of the decompiled program, but in a finite space: instead of keeping a full, unbounded execution stack, the decompiler collapses the stack into a finite context structure. That is, both the dynamic execution stack and the static context can be thought of as sequences of basic blocks, with the static context being a bounded sequence. [^4] The (dynamic) execution stack will include more data than continuations, such as values for expressions. However these are not relevant to our discussion, since the decompiler needs to cover all available code, regardless of run-time data values. Such data are also modeled statically in the decompiler, up to a (large) finite depth, as sets of values, instead of a single value in dynamic execution. These sets of values are kept per-context, much like in dynamic execution stack values appear in a specific stack depth, with specific contents of previous stack positions. The essence of the context-sensitivity algorithm is to decide which elements of the execution stack to keep at every point of modification, i.e., at every jump instruction. Different dynamic executions that have the same context (because their differing elements have been dropped by the context-sensitivity algorithm) will be treated the same, with the analysis computing all possible values for a variable, instead of just a single value.
@@
The example of Figure 4 helps explain the intuition behind shrinking context sensitivity. The example illustrates the effects of shrinking context sensitivity, contrasted with Elipmoc’s transactional context-sensitivity, on the private function context, for a series of block transitions. Blocks 0x1ca and 0x1b9 likely perform function calls as, following their execution, they leave the stack with continuations (0x1d3 and 0x1c3, respectively) pushed to it. For shrinking context sensitivity, when these continuations are reached, the most recent blocks, up to the block that pushed the continuation on the stack are dropped. This leaves room to maintain other information, within the same maximum context depth: shrinking context sensitivity reaches block 0x1d3 with the same calling context as the initial block 0x1ca, retaining crucial information about how the latter was reached. In contrast, transactional context-sensitivity reaches its maximum context depth and has to remove the oldest element from the stack, when it analyzes block 0x1d3. This will mean that if block 0x1ca was reachable under more than one calling context, upon reaching block 0x1d3, transactional context-sensitivity is unable to differentiate these contexts, merging them all into one.
@@
As a reminder, in predicate PrivateCallAndContinuation(caller: B, cont: B), the continuation does not have to be the block to return to after the call performed by block caller (although this is what it would be in a straightforward, unoptimized compilation of a simple call). It can instead be the return block for the caller’s caller (in case of tail calls), or the entry block of another called function (in case of chained calls), or any other block determined by complex optimizing compilation patterns.
@@
  • $\bullet$
    if a likely private function entry is found, push the caller block in the private part of the context, simultaneously dropping the oldest block in the context. Do the same if the transition is a likely function return, but it cannot be matched with a call in the context. Matching is performed by checking the continuation that the earlier likely call has pushed against the one the return is transitioning to;
  • $\bullet$

### v2

@@ Abstract.
We compare Shrnkr to state-of-the-art decompilers, both static-analysis- and symbolic-execution-based. In a standard benchmark set, Shrnkr scales to over 99.5% of contracts (compared to $\sim$95for Elipmoc), covers (i.e., reaches and manages to decompile) 67% more code than Heimdall-rs, and reduces key imprecision metrics by over 65%, compared again to Elipmoc.
Program Analysis, Smart Contracts, Decompilation, Datalog, Ethereum
copyright: noneccs: Theory of computation Program analysisccs: Software and its engineering General programming languagesccs: Security and privacy Software and application security
@@ 1. Introduction
From a technical standpoint, the problem of EVM decompilation is especially challenging. The EVM bytecode language is extremely low-level with respect to control flow, replacing all execution-control constructs with jump instructions to an address popped from the execution stack. That is, all control-flow statements (e.g., conditionals, loops, function calls, function returns) are translated into jumps to an address that is a run-time value of the low-level program. The challenge of EVM decompilation, thus, is to derive a higher-level representation, including functions, calls, returns, and structured control flow, from EVM bytecode. As a program analysis challenge, it has been the domain for applying several techniques. The primary distinction is between symbolic execution approaches (Kolinko and Palkeo, 2020; Becker, 2023) and static analysis (Grech et al., 2019, 2022) approaches. Symbolic-execution-based decompilers are easier to develop, naturally produce partial results (i.e., can always produce something, rather than failing on a smart contract in its entirety), yet are often vastly incomplete, failing to even discover a significant portion of the code. In contrast, static-analysis-based decompilers typically require much heavier development effort, cover (nearly) all code, but can fail to scale or can produce imprecise output.
@@
The technical essence of Shrnkr lies in several improvements over past static-analysis-based approaches: more precise and scalable static modeling, control-flow normalization via cloning, and pre-analysis-guided elimination of spurious calls. One key novelty is responsible for the lion’s share of the benefit: the fundamental static model is improved, by use of a new kind of static context kept inside the decompiler. That is, the decompiler maintains as its current control-flow history (i.e., how the execution got to the currently-analyzed statement) a list of basic blocks that is updated under a different algorithm. The new logic, dubbed shrinking context sensitivity, aggressively shrinks the context when a likely matching call-return or chained-call pattern is observed.
@@
2. Background
We next introduce the setting of this work: smart contracts, decompilation, context sensitivity.
2.1. EVM Smart Contracts
Smart contracts are small programs (typically up to around 1,000 lines of high-level code, limited to 24KB in binary form in Ethereum) stored on a persistent blockchain as part of its state. They are typically written in a high-level programming language, with Solidity being by far the most widely used. Solidity, which is the setting of our work, dominates over all other languages in terms of adoption. At the high-level, a smart contract defines a set of external/public functions, which are its public entry points through which Externally-Owned Accounts (EOAs) or other smart contracts can interact with the contract, and a set of persistent storage variables which are part of the contract’s state on the blockchain. Code reuse is facilitated through the use of internal (a.k.a. private) functions, inheritance, and library contracts. Solidity is a statically-typed language supporting operations on a number of value types (signed and unsigned integers, bytes, boolean), dynamic-length arrays, and associative mappings, as well as complex types combining the above.
@@ 2.1. EVM Smart Contracts
In addition, compiler version and settings greatly affect the produced bytecode. The release of Solidity v0.8.0 (Solidity Team, 2020) introduced checked arithmetic and employed v2 of the ABI encoder, greatly increasing the number of internal functions. Since Solidity v0.8.13 (Solidity Team, 2022), a new compilation pipeline became stable, with plans to make it the default in a future release (Solidity Team, 2024). This new Yul/viaIR pipeline involves Yul: a standardized, exportable intermediate language/representation (IR), which is also integrated into the Solidity language as inline assembly (Chaliasos et al., 2022). The Yul/viaIR pipeline enables deeper optimizations and more auditable code generation than the currently default “legacy” pipeline.
2.2. Context Sensitivity
Context sensitivity (Shivers, 1991; Milanova et al., 2005; Smaragdakis et al., 2011; Emami et al., 1994; Sharir and Pnueli, 1981) has a long history in static program analysis, with work in over 3 decades. Every static analysis that computes the flow of abstract values through the program is trying to approximate the solution to an undecidable problem. As a result, it faces challenges with respect to both scalability and precision. Addressing such challenges requires careful design decisions and context-sensitivity has offered a good way to balance these needs.
Context-sensitivity associates program variables (and sometimes heap objects) with context information and distinguishes the values not just on the basis of variables but on the basis of variable+context combinations. Analysis inferences for multiple executions that result in the same context will be merged, but stay differentiated from inferences associated with different contexts. Call-site-sensitivity, i.e. using one or more previous call-sites as context information, has seen success in analyzing functional languages (Shivers, 1991) and low-level imperative languages(Emami et al., 1994). For object-oriented languages, the use of the receiver object(s) as context information has been the state-of-the-art context sensitivity abstraction since its introduction. Section 7 offers more detailed pointers and comparison with past work.
2.3. EVM Decompilation
@@
We define the problem of EVM bytecode decompilation (a.k.a. binary lifting) as the derivation of high-level control-flow constructs and program structure from EVM bytecode. One can view the problem as the attempt to reconstruct a high-level program [^2] Notably, none of the EVM decompilers produce code that can be re-compiled. This does not diminish the value of EVM decompilation: the output of state-of-the-art decompilers is typically excellent both for human consumption and for writing automated program processing tools (e.g., static analyzers (Smaragdakis et al., 2021), symbolic-execution tools (Gritti et al., 2023; Ruaro et al., 2024), or program verification engines (Grossman et al., 2017; Albert et al., 2023, 2020)). from a low-level, stack-based intermediate representation (IR), where all control flow is represented in a continuation-passing style (CPS) form. For instance, a function call is done by pushing a continuation on the stack (the address of the basic block to return to), then pushing the function’s entry block address, and jumping. All control-flow patterns, such as in-function branching, tail calls, calls in-sequence, passing a return value of a call as an argument to another, etc., are represented as complex sequences of pushing continuations and eventually jumping to the first.
In the setting of EVM decompilation, the dynamic nature of the JUMP operations creates the need for whole-program reasoning, in order to compute a program’s control-flow graph. The Gigahorse/Elipmoc framework (Grech et al., 2019, 2022) has addressed this need by introducing a global context-sensitive control-flow graph/points-to analysis as the backbone of its decompiler.
This context-sensitive global control-flow graph is then used by the Elipmoc framework (Grech et al., 2022), which we extend, to identify potential call-sites, which are in turn used to compute function boundaries. Lastly, after the function boundaries are computed, their number of arguments and return arguments are inferred.
The Gigahorse lifter employed a N-call-site (or jump-site) context-sensitivity algorithm, while Elipmoc introduced a composite approach that included the identity of the public entry point and the 8 last call-sites that are likely private function calls or returns. The evaluation of the Elipmoc publication highlighted the key importance the context-sensitivity algorithm plays in the decompiler’s scalability and precision.
@@ 3.1. Compiler Translation
To glimpse the low-level complexity of compiled smart contracts, we consider a simple example. The program of Figure 1 contains two external functions that accept various parameters and perform fund-transfer operations. As we discuss later, the expression amt - defaultFee - feeA - feeB on line 5 actually performs three function calls to a private function used to check subtraction for underflow.
1interface IERC20 { function transfer(address to, uint256 value) external returns (bool); }
2contract DecomptTest {
@@
4 function transWFee(address tok, address to, uint256 amt, uint256 feeA, uint256 feeB) external
5 { IERC20(tok).transfer(to, amt - defaultFee - feeA - feeB); /* 3 private function calls */ }
6 function simpleTransfer(address tok, address to, uint256 amt) external
7 { IERC20(tok).transfer(to, amt); }
8}
@@ 3.2. Public Function Patterns
As public functions are not inherent in the EVM, high-level languages adhere to the contract Application Binary Interface (ABI) (Solidity, [n. d.]), which specifies how input and output data are to be encoded when interacting with a smart contract. Per the ABI, the first 4 bytes provided in a smart contract invocation are the function selector, used to identify the public function being called.
@@
0x25: PUSH2 0x38 // push function address of transWFee() public function
@@
Figure 2 shows the bytecode implementing the function selector logic for our example. These compiler-produced patterns implementing the function-selector logic have been used by past tools (Grech et al., 2019, 2022; Albert et al., 2018) to identify public function entries. However even detecting such simple patterns can have challenges. Looking at the previous code segment it is trivial for a local analysis to deduce that the EQ statement 0x24 operates on the function selector, since the input variable is loaded in the same block. EQ statement 0x2f also checks the function selector, however an inter-block analysis is required to be able to deduce this. The state-of-the-art Elipmoc binary lifter requires computing the public function entries before performing its global control-flow graph analysis (Grech et al., 2022, Figure 4). In order to achieve this, it uses an approximation based on a local-only analysis that does not verify that the selector is actually used.
In Section 5.2 we propose a 2-phase global analysis that, among others, tackles this problem.
@@ 3.3. Private Function Patterns
0x12d: DUP3 // position argument in stack
@@
0x10b: PUSH20 0xffffffffffffff...ffffffffffffffffff
0x121: AND // masks arg’s upper 12 bytes off
@@
(a) Simple Private Function Call
@@
0x60: PUSH2 0x72 // pushes cont. for 3rd call
@@
0x66: PUSH2 0x72 // pushes cont. for 2nd call
@@
0x6e: PUSH2 0x1c7 // push safeSub func address
@@
0x72: JUMPDEST // cont. address for 2nd, 3rd call
0x73: PUSH2 0x1c7 // push safeSub func address
@@
0x77: JUMPDEST ...
(b) Optimized Chained Private Function Calls
Figure 3. Private Function Call Patterns
The basic pattern for private function calls, presented in Figure 3(a), has been identified in past literature (Grech et al., 2019, 2022; He et al., 2023). A basic block makes a call to internal function cleanup_t_uint160 at offset 0x109, having first pushed the bytecode offset at which it wants to return after the called function’s execution completes. The return block of a function is a locally unresolved block that jumps back to the continuation block pushed by its caller.
In case of optimized code, a block performing an internal function call can also push the continuations of future calls. This pattern is typically produced in cases of calls that can be chained together (such as complex arithmetic expressions). As an example, consider the optimized compilation of expression ‘amt - defaultFee - feeA - feeB’ from our example program. The subtraction operations are actually function calls, to a checked-subtraction function, so the expression should be thought of as ‘safeSub(safeSub(safeSub(amt,defaultFee),feeA),feeB)’.
The resulting bytecode can be seen in Figure 3(b).
In this optimized case, block 0x58 will set the stack so that the 3 checked sub-operations are chained. Earlier work on the Elipmoc decompiler includes a function reconstruction algorithm (Grech et al., 2022, Figure 6) that recursively infers these chained function calls. It assumes that each low-level block will map to one high-level function call. However, in optimized code, the same low-level block can be used to perform more than one high-level function call. This can be seen in the above example, which pushes the address of block 0x72 on the stack twice, implementing in this way the last two checked-subtraction operations. Effectively, if we think of our high-level code as ‘safeSub1(safeSub2(safeSub3(amt,defaultFee),feeA),feeB)’ then safeSub2 and safeSub3 are a single low-level instruction, the same for both calls.
Such complex patterns evade past function reconstruction algorithms. In Section 5.1 we propose a cloning-based technique that identifies such low-level blocks with different high-level uses, and clones them, recovering precision of decompilation output.
@@ 3.4. Our Context
Elipmoc’s combination of scalability, precision, and completeness, paired with its expressive IR, have established it as a dominant lifter for EVM bytecode. Research tools for diverse program analysis applications have been implemented on top of Elipmoc. The applications include static-analysis (Brent et al., 2020; Lagouvardos et al., 2020; Kong et al., 2023; Liao et al., 2022; Smaragdakis et al., 2021; Grech et al., 2020; Munir and Reichenbach, 2023; Yang et al., 2024; Zhong et al., 2024; Zhang et al., 2025; Lagouvardos et al., 2025a), symbolic execution (Ruaro et al., 2024; Gritti et al., 2023), and deep learning (Zhao et al., 2023).
Shrnkr’s novel techniques mainly improve the decompiler’s context-sensitive global control-flow graph with the introduction of the shrinking context sensitivity variant described in Section 4 and its tuning via incompleteness in Section 5.2. Section 5.1 describes the introduction of a block cloning transformation step, which is performed before the global analyses and helps Shrnkr produce normalized decompilation output.
@@ 4. Shrinking Context-Sensitivity
Past work (Grech et al., 2022, 2019) has established a context-sensitive global control-flow-graph analysis as the backbone of a decompiler. That is, the decompiler abstractly simulates all possible executions of the decompiled program, but in a finite space: instead of keeping a full, unbounded execution stack, the decompiler collapses the stack into a finite context structure. That is, both the dynamic execution stack and the static context can be thought of as sequences of basic blocks, with the static context being a bounded sequence.
The essence of the context-sensitivity algorithm is to decide which elements of the execution stack to keep at every point of modification, i.e., at every jump instruction. Different dynamic executions that have the same context (because their differing elements have been dropped by the context-sensitivity algorithm) will be treated the same, with the analysis computing all possible values for a variable, instead of just a single value.
@@
The example of Figure 4 helps explain the intuition behind shrinking context sensitivity. The example illustrates the effects of shrinking context sensitivity, contrasted with Elipmoc’s transactional context-sensitivity, on the private function context, for a series of block transitions. Blocks 0x1ca and 0x1b9 likely perform function calls as, following their execution, they leave the stack with continuations (0x1d3 and 0x1c3, respectively) pushed to it. For shrinking context sensitivity, when these continuations are reached, the most recent blocks, up to the block that pushed the continuation on the stack, are dropped. This leaves room to maintain other information, within the same maximum context depth: shrinking context sensitivity reaches block 0x1d3 with the same calling context as the initial block 0x1ca, retaining crucial information about how the latter was reached. In contrast, transactional context-sensitivity reaches its maximum context depth and has to remove the oldest element from the stack, when it analyzes block 0x1d3. This will mean that if block 0x1ca was reachable under more than one calling context, upon reaching block 0x1d3, transactional context-sensitivity is unable to differentiate these contexts, merging them all into one.
@@
As a reminder, in predicate PrivateCallAndContinuation(caller: B, cont: B), the continuation does not have to be the block to return to after the call performed by block caller (as it would be in a straightforward, unoptimized compilation of a simple call). It can instead be the return block for the caller’s caller (in case of tail calls), or the entry block of another called function (in case of chained calls), or any other block determined by complex optimizing compilation patterns.
@@
  • if a likely private function entry is found, push the caller block in the private part of the context, simultaneously dropping the oldest block in the context. Do the same if the transition is a likely function return that cannot be matched with a call in the context. Matching is done by comparing the continuation (that the earlier likely call has pushed) with the one the return is going to;